# Draw2Cut: Direct On-Material Annotations for CNC Milling

### v1

@@ 3. Design Criteria
We propose using handwritten instructions (e.g., lines and symbols) to bridge the gap between virtual modeling and physical fabrication. To develop the design criteria, we first investigated the envisioned interactions to identify possible input and output methods. We subsequently developed a fast prototyping system using an existing tool to identify the technical requirements. This section focuses on the following questions.
@@ 3.1. Envisioned Interaction
  • The first is the use of colors to distinguish between the activities inspired by a laser cutter. When performing tasks that involve varying levels of power or speed, laser cutting employs a technique known as color-mapping (colormapping, ), which allows users to assign different settings to different regions of an image depending on the color.
@@
  • The third is the types of products that can be made by a three-axis CNC machine. The three-axis CNC machines, with the mill bit moving from top to bottom, are used for vertical subtraction. Consequently, they can be used to carve contour patterns and pockets or holes of various shapes, but only for limited 3D engraving (cut a part with flat features of varied depth) (2.5D, ). Based on these three types, we choose the three colors to cover all conceivable cutting results.
To summarize, we chose the three existing colors (purple, red, and green) because of their significant variances in HSV values, which improves the color identification accuracy. Pen colors were used to symbolize the various functions (Figure 2c). For example, a purple pen indicates a blueprint. The red and green pen symbols depict the tool path instructions. Symbols can be systematized to ensure rigidity, and people typically use them in their daily lives. Past research has explored the strong expressiveness of freehand annotation (song2006modelcraft, ), and certain symbols (such as crosses) have well-known meanings that allow users to draw instructions naturally. Inspired by the color mapping of laser cutting, the color language of Draw2Cut can be customized by the user.
@@
CAD is the dominant tool for planning complex 3D architectures with various elements, designing millimeter-accurate gears, and mass-producing goods for standard models. We do not intend to replace general-purpose CAD with Draw2Cut. The proposed system should have a specific scope for use in the targeted applications. Two promising applications are identified. The first is the creation of customized artifacts. This can be a one-of-a-kind artifact (zoran_freed_2013, ) recording improvisational thoughts (tian_matchsticks_2018, ) or handwritten identifications. Draw2Cut also reduces the skill level required for its use. Children can even create models quickly by drawing on wood. The second is the fabrication requirements for measuring, scanning, or constructing real items as references during the manufacturing process. Examples include joint customization and furniture modification.
@@ 4. Draw2Cut Workflow
Figure 6. Underlying technical process involving data collection, virtual space construction, image interpretation, and trajectory planning.
@@
After confirming the preview, the system updates the generated G-code, uploads it to the CNC system, and automatically moves to cut the output. After cutting is completed, the user can assemble, paint, or sand-cut pieces to complete the final product.
@@ 7.1. Basic Advantages
Figure 10. Edge joint (first row), and T-joint (second row) made by Draw2Cut
@@ 7.2. Applicability to Traditional Woodworking Tasks
It considers the cut depth, distance, and wood softness while creating modern curves and flexible shapes. We demonstrated how to create a ringed fence using a straight wood piece (Figure 12(e)). In addition to this demonstration, there are other projects that combine kerf manufacturing with high-detail engraving, usually to improve aesthetics or produce useful items with fine details, such as artists and wearable art (woodart, ).
@@ 7.3. Workshop
Figure 13. Products designed and fabricated in the workshop: (a) user’s drawing results; (b) Draw2Cut interpretation of drawing intent and visualization in red trajectories; (c) CNC-fabricated output; and (d) how users interact with the output.
@@ 8.2. Comparison between Draw2Cut, CAD, and Traditional Fabrication
Both CAD and Draw2Cut can lower the typical fabrication barriers for novice and experienced users in different aspects. Draw2Cut excels in offering high freedom and customization in manufacturing. Each artifact produced using Draw2Cut is unique, reflecting the designer’s individual style. By leveraging users’ sketching skills, Draw2Cut lowers the entry barrier for personal fabrication, enabling even children to design and manufacture toys. In contrast, manual woodworking requires specialized expertise, and CAD users need both hardware and programming skills. CAD offers clear advantages in handling complex designs. Its interface enables users to solve geometric problems with constraint-based relationships and design intricate interrelated components using timeline-based modeling. Additionally, CAD supports importing external libraries, such as joint libraries (tian_matchsticks_2018, ), perspective tools (larsson_tsugite_2020, ), or material composition analysis for parameter setting (dogan_sensicut_2021, ).
@@ 9.2. Hardware Limitations
We tried several methods to increase the accuracy and collected multi-frame data to filter out noise when creating a virtual workspace. Additionally, we discovered that the offset was stable; therefore, we manually adjusted for it. In the future, two methods can be used to increase the accuracy. First, we may test the system using a high-resolution depth camera. Second, numerous depth cameras may be positioned at various locations throughout the workspace. RGB cameras can be complemented by NIR cameras to detect hidden QR codes for more unobtrusive surface registration (dogan_infraredtags_2022, ; dogan_brightmarker_2023, ). Furthermore, CNC have an infinite number of end mills for various cuts. Each milling bit required its own feed rate and spindle speed. Throughout the development and experimentation, we continued to use one type (tapered ball nose carving end mill 1/16 in), which was narrow and had a slight slope. This mill bit is suitable for 3D carving (ballnose, ) but not for an assembly pocket. Because the upper half was thicker than the lower half, the sliced shape expanded when cut to a specific depth. In the future, we will generalize this language to cover multiple forms of mill bits.
@@ 9.3. Generalizability to Other Subtractive Manufacturing Processes
First, we examine the distinctions between other machines. Laser and waterjet cutting typically cut through materials that cannot be used for 3D carving jobs (dogan_structcode_2023, ). If the existing Draw2Cut language is used directly, only purple and red pens are required. The properties of the laser cutter are investigated in detail. When different powers are used, they can cut through materials (such as metal) and heat specific lines of plastic sheets, allowing them to be folded to perform origami activities (kusunoki2024integration, ). Consequently, new colors may be used to describe the power and speed of the laser cutter. We investigated the commercial use of waterjet and discovered that it has high requirements for the intricacy of 2D designs. Therefore, we should consider using different colors to represent the dashed cutting line or drill pattern. In addition to the current auto-smoothing function, the visualization interface requires a regularization function.
@@ Acknowledgements.
This work was supported by JST ACT-X Grant Number JPMJAX210P, JSPS KAKENHI Grant Number JP23K19994, JST AdCORP Grant Number JPMJKB2302 and a collaborative research fund between Mercari Inc. R4D and RIISE, University of Tokyo.

### v3

@@ 3. Design Criteria
We propose using handwritten instructions (e.g., lines and symbols) to bridge the gap between virtual modeling and physical fabrication. To develop the design criteria, we first investigated the envisioned interactions to identify possible input and output methods. We subsequently developed a fast prototyping system using an existing tool to identify technical requirements. This section focuses on the following questions.
@@ 3.1. Envisioned Interaction
  • The first is the use of colors to distinguish between the activities inspired by a laser cutter. When performing tasks that involve varying levels of power or speed, laser cutting employs a technique known as color-mapping (PPLD, 2024), which allows users to assign different settings to different regions of an image depending on the color.
@@
  • The third is the types of products that can be made by a three-axis CNC machine. The three-axis CNC machines, with the mill bit moving from top to bottom, are used for vertical subtraction. Consequently, they can be used to carve contour patterns and pockets or holes of various shapes, but only for limited 3D engraving (cut a part with flat features of varied depth) (CNC, 2024). Based on these three types, we choose the three colors to cover all conceivable cutting results.
To summarize, we chose the three existing colors (purple, red, and green) because of their significant variances in HSV values, which improves the color identification accuracy. Pen colors were used to symbolize the various functions (Figure 2c). For example, a purple pen indicates a blueprint. The red and green pen symbols depict the tool path instructions. Symbols can be systematized to ensure rigidity, and people typically use them in their daily lives. Past research has explored the strong expressiveness of freehand annotation (Song et al., 2006), and certain symbols (such as crosses) have well-known meanings that allow users to draw instructions naturally. Inspired by the color mapping of laser cutting, the color language of Draw2Cut can be customized by the user.
@@
CAD is the dominant tool for planning complex 3D architectures with various elements, designing millimeter-accurate gears, and mass-producing goods for standard models. We do not intend to replace general-purpose CAD with Draw2Cut. The proposed system should have a specific scope for use in the targeted applications. Two promising applications are identified. The first is the creation of customized artifacts. This can be a one-of-a-kind artifact (Zoran and Paradiso, 2013) recording improvisational thoughts (Tian et al., 2018) or handwritten identifications. Draw2Cut also reduces the skill level required for its use. Children can even create models quickly by drawing on wood. The second is the fabrication requirements for measuring, scanning, or constructing real items as references during the manufacturing process. Examples include joint customization and furniture modification.
@@ 4. Draw2Cut Workflow
Figure 6. Technical processes for data collection, virtual space construction, image interpretation, and trajectory planning.
@@
After confirming the preview, the system updates the generated G-code, uploads it to the CNC system, and automatically moves to cut the output. After cutting is complete, the user can assemble, paint, or sand-cut pieces to complete the final product.
@@ 7.1. Basic Advantages
Figure 10. Edge joint (first row), and T-joint (second row) made by Draw2Cut.
@@ 7.2. Applicability to Traditional Woodworking Tasks
It considers the cut depth, distance, and wood softness while creating modern curves and flexible shapes. We demonstrated how to create a ringed fence using a straight wood piece (Figure 12(e)). In addition to this demonstration, there are other projects that combine kerf manufacturing with high-detail engraving, usually to improve aesthetics or produce useful items with fine details, such as artists and wearable art (tortec, 2024).
@@ 7.3. Workshop
Figure 13. Products designed and fabricated in the workshop: (a) user’s drawing results, (b) Draw2Cut interpretation of drawing intent and visualization in red trajectories, (c) CNC-fabricated output; and (d) how users interact with the output.
@@ 8.2. Comparison between Draw2Cut, CAD, and Traditional Fabrication
Both CAD and Draw2Cut can lower the typical fabrication barriers for novice and experienced users in different aspects. Draw2Cut excels in offering high freedom and customization in manufacturing. Each artifact produced using Draw2Cut is unique, reflecting the designer’s individual style. By leveraging users’ sketching skills, Draw2Cut lowers the entry barrier for personal fabrication, enabling even children to design and manufacture toys. In contrast, manual woodworking requires specialized expertise, and CAD users need both hardware and programming skills. CAD offers clear advantages in handling complex designs. Its interface enables users to solve geometric problems with constraint-based relationships and design intricate interrelated components using timeline-based modeling. Additionally, CAD supports importing external libraries, such as joint libraries (Tian et al., 2018), perspective tools (Larsson et al., 2020), or material composition analysis for parameter setting (Dogan et al., 2021).
@@ 9.2. Hardware Limitations
We tried several methods to increase the accuracy and collected multi-frame data to filter out noise when creating a virtual workspace. Additionally, we discovered that the offset was stable; therefore, we manually adjusted for it. In the future, two methods can be used to increase the accuracy. First, we may test the system using a high-resolution depth camera. Second, numerous depth cameras may be positioned at various locations throughout the workspace. RGB cameras can be complemented by NIR cameras to detect hidden QR codes for more unobtrusive surface registration (Dogan et al., 2022b, 2023b). Furthermore, CNC have an infinite number of end mills for various cuts. Each milling bit required its own feed rate and spindle speed. Throughout the development and experimentation, we continued to use one type (tapered ball nose carving end mill 1/16 in), which was narrow and had a slight slope. This mill bit is suitable for 3D carving (Lieurance, 2016) but not for an assembly pocket. Because the upper half was thicker than the lower half, the sliced shape expanded when cut to a specific depth. In the future, we will generalize this language to cover multiple forms of mill bits.
@@ 9.3. Generalizability to Other Subtractive Manufacturing Processes
First, we examine the distinctions between other machines. Laser and waterjet cutting typically cut through materials that cannot be used for 3D carving jobs (Dogan et al., 2023a). If the existing Draw2Cut language is used directly, only purple and red pens are required. The properties of the laser cutter are investigated in detail. When different powers are used, they can cut through materials (such as metal) and heat specific lines of plastic sheets, allowing them to be folded to perform origami activities (Kusunoki et al., 2024). Consequently, new colors may be used to describe the power and speed of the laser cutter. We investigated the commercial use of waterjet and discovered that it has high requirements for the intricacy of 2D designs. Therefore, we should consider using different colors to represent the dashed cutting line or drill pattern. In addition to the current auto-smoothing function, the visualization interface requires a regularization function.
@@ Acknowledgements.
This work was supported by JST ACT-X Grant Number JPMJAX210P, JSPS KAKENHI Grant Number JP23K19994, JST AdCORP Grant Number JPMJKB2302 and a collaborative research fund between Mercari Inc. R4D and RIISE, The University of Tokyo.